The investor community is worried that Argentina will default on its loans (that is, that Argentina does not pay the interest promised on the bonds it has issued). Assume that Argentinean debt can be modeled as a 50-year stream of interest payments of 3 billion dollars each, taking place every 6 months, plus a final payment of 30 billion dollars in 50 vears time 12. a. Suppose that the annual effective required discount rate on Argentinas debt is 10.25%. How much does Argentina owe to its creditors? b. The Finance minister has recently proposed a debt restructuring plan according to which Argentina would like to see the interest payments reduced by 75%. That is. Argentina it would pay only 25 cents for each dollar in promised interest. If Argentina carries out this plan, how much money would the bondholders lose? c. One financial analyst commented that the minister actually wants to see the present value of the debt reduced by 75%. If this statement is true, are the creditors better or worse off than the original plan proposed in b.? Please explain why. [Note: you can answer this question even if you did not solve part b.]
Answer #1

(1) (a) Interest Amounts = $ 3 billion and Terminal Balloon Payment = $ 30 billion, Effective Interest Rate = 10.25 % and Tenure of Repayment = 50 years or 100 half-years

Semi-Annual Discount Rate = (10.25 / 2) = 5.125 %

Therefore, Present Value of Repayment's = Amount Owed By Argentina = 3 x (1/0.05125) x [1-{1/(1.05125)^(100)}] + 30 / (1.05125)^(100) = $ 58,3439 billion ~ $ 58.34 billion

(b) if the interest payments are reduced by 75 %, the new interest payments = 3 x (1-0.75) = $ 0.75 billion

Debt Repaid = 0.75 x (1/0.05125) x [1-{1/(1.05125)^(100)}] + 30 / (1.05125)^(100) = $ 14.738 billion ~ $ 14.74 billion

Amount Lost by Bondholders = 58.34 - 14.74 = $ 43.6 billion

(c) If the present value of the debt is reduced by 75 %, the new debt value = 58.34 x (1-0.75) = $ 14.585 billion

As is observable, the investors are worse off in this case as compared to the previous strategy of reducing interest payments. This happens because for a long tenure such as this, a major part of the present value is the large number of interest repayments which even though reduced by the discounting factor form a high value owing to the number of interest repayments.